If a pice of wood is 5.5 cm and I get 5.62cm what is the percent error
1
Expert's answer
2020-12-16T20:18:02-0500

"To \ determine \ the \ error, \ firstly \ we \ subtarct \ the \ wood \ gotten \ and \ the \ wood \ available: \"

"5.62 - 5.5 = 0.12 (sm)."

"Then \ need \ to \ find \ error \ in \ percentage \ with \ respect \ to \ the \ wood \ available: \"

"\frac{0.12}{5.5} = 0.0218 = 2.18\ \%"